OpenStudy (anonymous):

f(x) = (2x+1)^(1/2) g is the inverse of f(x) g'(3) = ? Shouldn't it be 3/2?

OpenStudy (ranga):

Can you show the steps you took to find the inverse function?

OpenStudy (anonymous):

(2x + 1)^(1/2) = 3 Therefore, x = 4 Now, g'(3) = 1/[2/(2(4)+1)^(1/2)] = 3/2